# Spectral selectors on lens spaces and applications to the geometry of the group of contactomorphisms

### v1

@@ 5. A spectral pseudonorm
$\nu_{\ast}\big{(}\Pi(\widetilde{\phi})\big{)}\leq\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}\leq 2\pi+\frac{2\pi}{k}$ (21)
for every $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. Using periodicity of the spectral selectors (Theorem 1.1 (vi)) and the fact that the sequence of spectral selectors $c_{j}$ is non-decreasing we have $c_{+}(\widetilde{\phi})\leq c_{-}(\widetilde{\phi})+2\pi$, which implies the second inequality in (21). For the first inequality, it is enough to find $N\in\frac{2\pi}{k}\cdot\mathbb{Z}$ such that $\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}$. Suppose first that $\nu(\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}$, and pose $N=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}$. By Theorem 1.1 (v) we then have
$\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{r_{-N}}\cdot\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}\,,\,-\left\lfloor c_{-}(\widetilde{r_{-N}}\cdot\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}\,\right\}$
@@
Similarly, if $\nu(\widetilde{\phi})=\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}=:N$ then $\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}}$. ∎
@@
It follows from [4, Corollary 4.12] that on the universal cover of the identity component of the contactomorphism group of the unit cotangent bundle of the torus $\mathbb{T}^{n}$ for $n\geq 2$ the difference of the invariants $c_{+}$ and $c_{-}$ defined in [4] is unbounded. This difference with respect to (21) might be related to the fact that the identity component of the contactomorphism group of the unit cotangent bundle of the torus does not contain positive loops. It would be interesting to investigate if on the other hand the difference of the invariants $c_{+}$ and $c_{-}$ of [4] is bounded on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. This would then imply as in Section 5 that the induced norm on $\operatorname{Cont}_{0}(L_{k}^{2n-1},\xi_{0})$ is bounded, and therefore answer partially a question in [12, Example 2.21].
@@
is a norm. Moreover, $\nu^{\prime}$ is invariant by conjugation if and only if so is $\nu$. This trick (which is similar to one used in [7]) can be applied to our pseudonorm $\nu$, with $c=\frac{2\pi}{k}$, to obtain a stably unbounded conjugation invariant norm $\nu^{\prime}$ on $\widetilde{\operatorname{Cont}_{0}}(L_{k}^{2n-1},\xi_{0})$. Since $\nu$ takes values in $\frac{2\pi}{k}\cdot\mathbb{Z}$, if $\nu$ is already a norm then $\nu^{\prime}\equiv\nu$. Section 5 holds also for $\nu^{\prime}$, indeed for any element $\widetilde{\phi}\neq\widetilde{\mathrm{id}}$ we have
$\nu^{\prime}(\widetilde{\phi})=\max\left\{\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{\frac{2\pi}{k}},-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{\frac{2\pi}{k}},\frac{2\pi}{k}\right\}\leq\frac{2\pi}{k}\cdot\max\big{\{}\nu_{\operatorname{osc}}(\widetilde{\phi}),1\big{\}}=\frac{2\pi}{k}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$

### v2

@@ 5. A spectral pseudonorm
$\nu_{\ast}\big{(}\Pi(\widetilde{\phi})\big{)}\leq\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}\leq 2\pi+T_{\underline{w}}$ (21)
for every $\widetilde{\phi}\in\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Using periodicity of the spectral selectors (Theorem 1.1 (vi)) and the fact that the sequence of spectral selectors $c_{j}$ is non-decreasing we have $c_{+}(\widetilde{\phi})\leq c_{-}(\widetilde{\phi})+2\pi$, which implies the second inequality in (21). For the first inequality, it is enough to find $N\in T_{\underline{w}}\cdot\mathbb{Z}$ such that $\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}$. Suppose first that $\nu(\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}$, and pose $N=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}$. By Theorem 1.1 (v) we then have
$\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\max\left\{\,\left\lceil c_{+}(\widetilde{r_{-N}}\cdot\widetilde{\phi})\right\rceil_{T_{\underline{w}}}\,,\,-\left\lfloor c_{-}(\widetilde{r_{-N}}\cdot\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}\,\right\}$
@@
Similarly, if $\nu(\widetilde{\phi})=-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}$ then, posing $N=\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}$ we have
$\nu(\widetilde{r_{-N}}\cdot\widetilde{\phi})=\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}}-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}}\,.$
@@
It follows from [4, Corollary 4.12] that on the universal cover of the identity component of the contactomorphism group of the unit cotangent bundle of the torus $\mathbb{T}^{n}$ for $n\geq 2$ the difference of the invariants $c_{+}$ and $c_{-}$ defined in [4] is unbounded. This difference with respect to (21) might be related to the fact that the identity component of the contactomorphism group of the unit cotangent bundle of the torus does not contain positive loops. It would be interesting to investigate if on the other hand the difference of the invariants $c_{+}$ and $c_{-}$ of [4] is bounded on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. This would then imply as in Section 5 that the induced norm on $\operatorname{Cont}_{0}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$ is bounded, and therefore answer partially a question in [13, Example 2.21].
@@
is a norm. Moreover, $\nu^{\prime}$ is invariant by conjugation if and only if so is $\nu$. This trick (which is similar to one used in [8]) can be applied to our pseudonorm $\nu$, with $c=T_{\underline{w}}$, to obtain a stably unbounded conjugation invariant norm $\nu^{\prime}$ on $\widetilde{\operatorname{Cont}_{0}}\big{(}L_{k}^{2n-1}(\underline{w}),\xi_{0}\big{)}$. Since $\nu$ takes values in $T_{\underline{w}}\cdot\mathbb{Z}$, if $\nu$ is already a norm then $\nu^{\prime}\equiv\nu$. Section 5 holds also for $\nu^{\prime}$, indeed for any element $\widetilde{\phi}\neq\widetilde{\mathrm{id}}$ we have
$\nu^{\prime}(\widetilde{\phi})=\max\left\{\left\lceil c_{+}(\widetilde{\phi})\right\rceil_{T_{\underline{w}}},-\left\lfloor c_{-}(\widetilde{\phi})\right\rfloor_{T_{\underline{w}}},T_{\underline{w}}\right\}\leq T_{\underline{w}}\cdot\max\big{\{}\nu_{\operatorname{osc}}(\widetilde{\phi}),1\big{\}}=T_{\underline{w}}\cdot\nu_{\operatorname{osc}}(\widetilde{\phi})\,.$